OpenStudy (anonymous):

Simplify 6xy(3xy)^-2 Answers: a.2xy b.2 divided by x y c.2 divided by the quantity three x y d.3xy

OpenStudy (anonymous):

same as\[\frac{ 6xy }{ 3xy ^{2} }\]

OpenStudy (anonymous):

6/3=2 and xy/xy^2 is 1/xy

OpenStudy (anonymous):

so the solutions is a?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2 times 1/xy= 2/xy

OpenStudy (anonymous):

So its B

OpenStudy (anonymous):

6xy*(3xy)^-2=6xy*9^-1*x^-2*y^-2 =3/ 2xy so the answer is c

OpenStudy (anonymous):

Ohhh ok than you!

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

yes, the answer is c.